OpenStudy (anonymous):

What are the foci of (x+3)^2/169 + (y+5)^2/144=1? (–8, 5) and (2, 5) (–2, 5) and (8, 5) (–8, –5) and (2, –5) (–2, –5) and (8, –5)

OpenStudy (anonymous):

a^2=169 b^2=144 c^2=a^2-b^2=25 so c=5. the foci is \(\pm 5 \) units left/right of the center of the ellipse. can you identify the center to get the coordinates of your foci?

OpenStudy (anonymous):

so which answer would it be?

OpenStudy (anonymous):

can you identify the center of the ellipse?

OpenStudy (anonymous):

(-3,-5)?

OpenStudy (anonymous):

so the foci is (-3+5, -5) and (-3-5, -5)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw...:)

OpenStudy (anonymous):

do you know this one? What is the focus of the parabola x =1/6 (y + 3)2 – 6 ?

OpenStudy (anonymous):

yes... get that equation into this form: 4a(x-h) = (y-k)^2 your vertex is (h,k) and the focus is "a" units away from the vertex... so you'll need to find "a"

OpenStudy (anonymous):

I`m not sure what the vertex is..

OpenStudy (anonymous):

ok... let's start with the equation... add 6 to both sides... write down what you have...

OpenStudy (anonymous):

oh i just realized that its supposed to be x=1/16 not just 6!